$SPAD/src/input d01fcf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
e:= (4.0*X[1]*X[3]*X[3]*exp(2.0*X[1]*X[3])/((1.0+X[2]+X[4])**2))::EXPR FLOAT
f:ASP4(FUNCTN):=retract e
ndim:=4
a:Matrix SF:=[[0.0,0.0,0.0,0.0]]
b:Matrix SF:=[[1.0,1.0,1.0,1.0]]
alpha:=2^ndim+2*ndim^2+2*ndim+1
minpts:=1000
maxpts:=5700
eps:=0.0001
lenwrk:=606
result:=d01fcf(ndim,a,b,maxpts,eps,lenwrk,minpts,-1,f)

———-

References^{[1] nothing}

4